# FlatFusion: Delving into Details of Sparse Transformer-based Camera-LiDAR Fusion for Autonomous Driving

### v1

@@ 3.3 Fusion & Projection
3.4 Partition
Table 5 gives performance comparisons, leading to the following conclusions:
Table 6: Selection of Group Size
@@ 3.4 Partition
Group Size: One key hyper-parameter of sparse windowed attention is the group size, i.e., the number of tokens per set to conduct attention. We investigate the performance of group size under 3 scales: 40, 80, 160, as in Table 6. We find that 80 is suitable, which is quite different from the conclusion in PTv3 [62], where they find larger group size up to 1024 could consistently bring performance gains. The discrepancy might be because they mainly focus on indoor point clouds [8], which are much denser compared to outdoor ones [2]. We visualize the partitioning results of different group size in Fig. 6. We can find that when using small group size, the receptive field is limited while large group would leads to bad locality, indicating the necessity of tuning.
3.5 Transformer Structure
Table 7: Design of Transformer Structure
@@ 3.5 Transformer Structure
Transformer [53] consists of attention, FFN, residual link, LayerNorm, and position encoding (PE). Following LLaMA family [52], we adopt FlashAttention [9] and SwiGLU [49]. Existing works like FlatFormer [43] and UniTR [57] both adopt PostNorm [53] while modern architecture [52] usually adopts PreNorm [4]. UniTR proposes to use an extra block-wise residual and post-norm. As for PE, we examine window-level local indices [57] or 3D position encoding.In Table 7, we find that PreNorm & 3D PE achieves best performance, aligned with modern findings [52].
3.6 Final Model
The final model could be summarized as: given images from surrounding cameras and point cloud from LiDAR as inputs, we use ResNet18 [13] followed by FPN [39] and Dynamic VFE [74] followed by 4 Flatten Window Transformer layers [62] as tokenizer to obtain $N_{C}*H*W$ image tokens and $N_{L}$ voxel tokens respectively. Then, we project LiDAR tokens to the pixel space and feed the mixed image and LiDAR tokens into 4 Flatten Window Transformer layers for 3D-to-2D fusion. After that, we project image tokens to 3D space by Partial Projection and use 4 Flatten Window Transformer layers for 2D-to-3D fusion. Finally, we densify [43, 57] the LiDAR feature followed by several convs and feed it into the 3D detection head [1]. Similar to [57], optionally, the image features could be projected by LSS [47, 17] to BEV and concatenated with the densified LiDAR features, which leads to better performance in the cost of more latency. The LiDAR feature is downsampled in z-axis with attentive pooling [56] between stages. All Transformer layers are PreNorm and use 3D PE with a group size 80. Please refer to appendix for more details about each component.
@@ 3.6 Final Model
Table 8: Performance of 3D detection on nuScenes validation and test set. FPS is measured on A100 GPU by default with cached intrinsic and extrinsic. : FPS on RTX3090. : FPS on RTX4090.
Method Present at Fusion Strategy Val Set Test Set FPS
NDS mAP NDS mAP
FUTR3D [5] CVPR’23 Instance Query 68.3 64.5 - - 3.3
UVTR [33] NeurIPS’22 2D-to-3D 70.2 65.4 71.1 67.1 1.8
BEVFusion [38] NeurIPS’22 BEV Concat 71.0 67.9 71.8 69.2 0.8
TransFusion [1] CVPR’22 LiDAR Query 71.3 67.5 71.6 68.9 6.1
BEVFusion [42] ICRA’23 BEV Concat 71.4 68.5 72.9 70.2 7.7
MSMDFusion [26] CVPR’23 Dual Space 72.1 69.3 74.0 71.5 2.1
ObjectFusion [3] ICCV’23 Instance Query 72.3 69.8 73.3 71.0 3.6
DeepInteraction [67] NeurIPS’22 Dual Attention 72.6 69.9 73.4 70.8 1.8
SparseFusion [64] ICCV’23 Instance Query 72.8 70.4 73.8 72.0 5.6
CMT [65] ICCV’23 Instance Query 72.9 70.3 74.1 72.0 6.4
UniTR [57] ICCV’23 Dual Attention 73.1 70.0 74.1 70.5 11.3 (10.3)
UniTR w/ LSS [57] ICCV’23 Dual Attention 73.3 70.5 74.5 70.9 9.3 (8.7)
FlatFusion - Dual Attention 73.4 70.2 74.3 70.6 11.0
FlatFusion w/ LSS - Dual Attention 73.7 70.6 74.7 71.1 10.1
4 Experiments
4.1 Setup
@@
Dataset: To evaluate the performance, we utilize nuScenes [2] , a large-scale public dataset specifically designed for autonomous driving applications. It includes 1,000 scenes, each 20 seconds long, with six cameras and one LiDAR sensor. We use the official train-val-test split (750-150-150). This multi-sensor setup captures a wide array of annotations, including 23 object classes and 8 attributes, over 1.4 million annotated 3D bounding boxes.
@@
4.2 Results
We give results of 3D objection Detection on nuScenes validation and test set in Table 8. We could observe that Flatfusion achieves strong performance yet with a high FPS, demonstrating the efficacy of the explored designs. Notably, compared to latest sparse Transformer based fusion methods, i.e., CMT [65], SparseFusion [64], UniTR [57], Flatfusion demonstrate explicit performance advantages.
Figure 7: Latency Analysis of Flatfusion. Latency is measured in nuScenes val set on RTX4090.
Table 9: Influence of Network Depth.
@@ 4.2 Results
(a) Block Depth
[…]
(b) BEV Bakcbone
4.3 Influence of Network Depth
In Table 9(a), we study the influence of network depth. As for the number of Transformer layers per block (LiDAR backbone, 3D-to-2D, 2D-to-3D), we find that 2 layers leads to underfit while 8 layers causes unstable training. As for BEV Backbones, we find that the commonly used SECOND [66] performas worse than customized ResNet [57], indicating the efficacy of residual connection.
4.4 Latency Analysis
In Fig. 7, we give the latency of each component in Flatfusion. We could observe that even though heavily optimized, the partition time is still one of the biggest overhead, which worth further study. Besides, the task head: densifying, BEV Bakcbone, and Transfusuion might be further light-weighted.
5 Conclusion
In this work, we delve into the details of sparse Transformer based camera-LiDAR fusion framework. We analyze the issues within existing methods and propose corresponding solutions. The proposed Flatfusion achieves state-of-the-art performance with a high FPS.
Appendix A Implementation Details
Formally, the input of the neural network is 6 images with the size of $256\times 704$ and the point cloud. The 6 images is first fed into ResNet18 [13] and FPN [39] to generate ($N_{c}\times H\times W$) image tokens where $N_{c}=6,H=256/8=32,W=704/8=88$. The point cloud is processed by Dynamic VFE [74] into $N_{L}$ tokens where $N_{L}$ is the number of non-empty voxels within the scene. Regarding the Flatten Window Transformer, we follow the design in [43]. Please refer to their paper [43] for more details. We give a brief description here to be self-contained:
Partition: With a point cloud $\{(x,y,z)\}$, they are first quantized by
$\bigl{(}\,\underbrace{\lfloor x/w_{x}\rfloor,\lfloor y/w_{y}\rfloor,\lfloor z/w_{z}\rfloor}_{\text{window coordinates}},\,\,\underbrace{x-\lfloor x/w_{x}\rfloor\cdot w_{x},y-\lfloor y/w_{y}\rfloor\cdot w_{y},z-\lfloor z/w_{z}\rfloor\cdot w_{z}}_{\text{local coordinates within window}}\,\bigr{)},$ (1)
where $(w_{x},w_{y},w_{z})$ is the window shape. Next, it sorts all points first by window coordinates and then by local coordinates within the window. This step turns the unordered point cloud into an ordered one, where points within the same window will be next to each other and then it partitions the point cloud into groups of equal sizes based on the sorted index.
Shift: Similar to Swin [41], to ensure tokens could "see" neighbors from all direction, it could either use $x$-major or $y$-major sorting. Besides, it could also shift the window $(w_{x}/2,w_{y}/2)$. Thus, there is 4 types of partitions: $x$, $x$-shift, $y$, $y$-shift and thus we have 4 Transformer layers per stage corresponding to these 4 partitions.
We feed tokens after Dynamic VFE into 4 Flatten Window layers to obtain the LiDAR tokens. Then, for 3D-to-2D fusion, LiDAR tokens are projected to pixel space according to their 3D coordinate. For each image, we feed its image tokens as well LiDAR tokens projected on this image into 4 Flatten Window layers, partitioned based on pixel coordinate. Next, for 2D-to-3D fusion, we project all image tokens with known depth, i.e., having point projected on, into 3D space, based on the depth. We feed image and LiDAR tokens into 4 Flatten Window layers, partitioned based on 3D coordinate. Finally, LiDAR tokens are densified with empty voxels padded with zeros and then fed into several 2D convs. Transfusion head [1] is used to output the 3D bounding boxes. Note that LiDAR tokens are downsampled gradually in $z$-axis with attentive pooling [56], where the value of $z$ in LiDAR tokenizer, 3D-to-2D fusion, 2D-to-3D fusion, Transfusion head is {32, 8, 2, 1} respectively. Optionally, we could project image tokens into BEV space by LSS [47] and concatenate with densified LiDAR tokens, before fed into Transfusion head.

### v2

@@ III-E Transformer Structure
TABLE IV: Design of partition
@@
Algorithm Dynamic Set 73.4 70.1 28.5ms
Flatten Window 73.4 70.2 17.3ms
Space-Fill Curves 73.1 70.0 22.6ms
Group Size 40 72.7 69.7 25.3ms
80 73.4 70.2 22.7ms
160 72.8 69.6 23.9ms
TABLE V: Design of transformer structure
@@
PostNorm 73.0 69.8 48.1ms
Block Residual [65] 72.6 69.3 49.2ms
Window-Level PE [65] 72.5 69.3 52.0ms
PreNorm + 3D PE 73.4 70.2 48.8ms
@@
TABLE VI: Performance of 3D detection on nuScenes validation and test set. FPS is measured on A100 GPU by default with cached intrinsic and extrinsic. : FPS on RTX3090. : FPS on RTX4090.
Method Present at Camera Branch Fusion Strategy NDS (val) mAP (val) NDS (test) mAP (test) FPS
TransFusion [1] CVPR’22 ResNet-50 [15] Instance Query 71.3 67.5 71.6 68.9 6.1
UVTR [39] NeurIPS’22 ResNet-101 [15] BEV Concat 70.2 65.4 71.1 67.1 1.8
BEVFusion [45] NeurIPS’22 Dual-Swin-T [44] BEV Concat 71.0 67.9 71.8 69.2 0.8
DeepInteraction [80] NeurIPS’22 ResNet-50 [15] Dual Attention 72.6 69.9 73.4 70.8 1.8
FUTR3D [5] CVPR’23 ResNet-101 [15] Instance Query 68.3 64.5 - - 3.3
MSMDFusion [32] CVPR’23 ResNet-50 [15] BEV Concat 72.1 69.3 74.0 71.5 2.1
BEVFusion [49] ICRA’23 Swin-T [48] BEV Concat 71.4 68.5 72.9 70.2 7.7
ObjectFusion [3] ICCV’23 Swin-T [48] Instance Query 72.3 69.8 73.3 71.0 3.6
SparseFusion [73] ICCV’23 ResNet-50 [15] Instance Query 72.8 70.4 73.8 72.0 5.6
CMT [74] ICCV’23 VoVNet-99 [34] Instance Query 72.9 70.3 74.1 72.0 6.4
FocalFormer3D  [6] ICCV’23 ResNet-50 [15] BEV Concat 73.1 70.5 73.9 71.6 -
UniTR  [65] ICCV’23 DSVT [64] Dual Attention 73.3 70.5 74.5 70.9 9.3 (8.7)
FSF [40] TPAMI’24 ResNet-50 [15] Instance Query 72.7 70.4 74.0 70.6 7.0
GraphBEV [57] ECCV’24 Swin-T [48] BEV Concat 72.9 70.1 73.6 71.7 7.1
FlatFusion - ResNet-18 [15] Dual Attention 73.7 70.6 74.7 71.1 10.1
III-F Final Model
The final model could be summarized as: given images from surrounding cameras and point cloud from LiDAR as inputs, we use ResNet18 [15] followed by FPN [46] and Dynamic VFE [92] followed by Flatten Window Transformer layers [71] as tokenizer to obtain camera tokens with a shape of $N_{C}*C$ and LiDAR tokens with a shape of $N_{L}*C$ respectively. Then, we project LiDAR tokens to the pixel space and feed the mixed image and LiDAR tokens into 4 Flatten Window Transformer layers for 3D-to-2D fusion. After that, we project image tokens to 3D space by Partial Projection and use 4 Flatten Window Transformer layers for 2D-to-3D fusion. Finally, we densify [50, 65] the LiDAR feature followed by several convs and feed it into the 3D detection head [1]. Similar to [65], optionally, the image features could be projected by LSS [54, 19] to BEV and concatenated with the densified LiDAR features, which leads to better performance in the cost of more latency. The LiDAR feature is downsampled in $z$-axis with attentive pooling [64] between stages. All Transformer layers are PreNorm and use 3D PE with a group size 80.
IV Experiments
IV-A Setup
@@
Dataset: We leverage nuScenes [2], a comprehensive public dataset meticulously curated for autonomous driving applications, to assess our performance. It includes 1,000 scenes, each 20 seconds long, with six cameras and one LiDAR sensor. We use the official train-val-test split (750-150-150). This multi-sensor setup captures a wide array of annotations, including 23 object classes and 8 attributes, over 1.4 million annotated 3D bounding boxes.
@@
Figure 7: Latency analysis for FlatFusion on the nuScenes validation set, measured with an RTX 4090.
IV-B Results
We give results of 3D objection detection on nuScenes validation and test set in Table VI. We could observe that FlatFusion achieves strong performance yet with a high FPS, demonstrating the efficacy of the explored designs. Notably, compared to the latest sparse Transformer based fusion methods, i.e., CMT [74], SparseFusion [73], UniTR [65], FlatFusion demonstrate explicit performance advantages.
IV-C Influence of Network Depth
In Table VII, we study the influence of network depth. As for the number of Transformer layers per block (LiDAR backbone, 3D-to-2D, 2D-to-3D), we find that 2 layers leads to underfit while 8 layers causes unstable training. Regarding BEV encoder, we find that the commonly used SECOND [75] performs worse than customized ResNet [65], indicating the efficacy of residual connection.
TABLE VII: Influence of network depth
@@ IV-C Influence of Network Depth
Block Depth 2 72.4 69.0 39.8
4 73.4 70.2 48.2
8 72.6 69.6 58.1
BEVEncoder SECOND (5+5) [49] 72.5 69.3 6.6
SECOND (8+8) [19] 72.8 69.3 9.2
ResNet14 [65] 73.4 70.2 8.7
IV-D Latency Analysis
In Fig. 7, we give the latency of each component in FlatFusion. We could observe that even though heavily optimized, the partition time is still one of the biggest overhead, which worth further study. Besides, the BEV backbone, as well as the densification and detection head, could potentially be further light-weighted.
V Conclusion
In this work, we delve into the details of sparse Transformer based camera-LiDAR fusion framework. We analyze the issues within existing methods and propose corresponding solutions. The proposed FlatFusion achieves state-of-the-art performance with a high FPS. We aspire for FlatFusion to stimulate future research endeavors centered around the design of efficient and precise multi-sensor fusion transformers.